OpenStudy (sleepyjess):

Write the expression in standard form. \(\Large\dfrac{7}{4-14i}\)

hartnn (hartnn):

multiply numerator and denominator by conjugate of 4-14i

OpenStudy (sleepyjess):

\(\Large -\dfrac{7}{53}+\dfrac{49}{106}i\) \(\Large\dfrac{7}{53}-\dfrac{49}{106}i\) \(\Large\dfrac{7}{53}+\dfrac{49}{106}i\) \(\Large -\dfrac{7}{53}-\dfrac{49}{106}i\)

hartnn (hartnn):

conjugate of 4-14i is ?

OpenStudy (sleepyjess):

4+14i

OpenStudy (sleepyjess):

i came up with \(\dfrac{28+98i}{212}\)

hartnn (hartnn):

thats correct :)

hartnn (hartnn):

now separate out the numerator

OpenStudy (sleepyjess):

\(\dfrac{28}{212}+\dfrac{98i}{212}\)?

hartnn (hartnn):

yep, now reduce those fractions

hartnn (hartnn):

anyways since both the parts are positive, you can guess the answer :)

OpenStudy (sleepyjess):

c?

OpenStudy (sleepyjess):

can you help me on like 2 more?

hartnn (hartnn):

yes, and sure :)